OpenStudy (lgbasallote):

Given: at 11:30 pm, temperature = 94.6 F at 12:30 am; temperature = 93.4 F room temperature (constant) = 70 F at what time is temperature = 98.6 F?

OpenStudy (lgbasallote):

\[\ln (\frac{93.4 - 70}{94.6-70}) = k\] \[\ln (0.9512) = k\] \[k = -0.0500\]

OpenStudy (lgbasallote):

\[\ln (\frac{94.6 - 70}{98.6 - 70}) = -0.0500t\] \[\ln (0.8601) = -0.0500t\] \[-0.1507 = -0.0500t\] \[t = 3.014 hours\]

OpenStudy (lgbasallote):

okay my question is actually just simple...

OpenStudy (lgbasallote):

how come 11:30 pm minus 3.014 hours is 8:29 pm?? or am i wrong???

OpenStudy (asnaseer):

3.014 hours is approximately 3 hours and 1 minute

OpenStudy (asnaseer):

convert the decimal part of the hours to minutes by multiplying by 60: 0.014 * 60 = 0.84

OpenStudy (lgbasallote):

then i did 0.84 x 60...

OpenStudy (lgbasallote):

it's 50 seconds

OpenStudy (lgbasallote):

so i approximate it to be 3 hours 1 minute? @asnaseer

OpenStudy (asnaseer):

yes

OpenStudy (lgbasallote):

thanks

OpenStudy (asnaseer):

yw :)

OpenStudy (lgbasallote):

lol so many people ran away when they saw differential equations..i just had a simple question

OpenStudy (asnaseer):

:)